# Frobenius intertwiners for $q$-difference equations

### v1

@@ 3. $p$adic $q$-hypergeometric functions
We have the following congruences for the normalized K-theoretic vertex functions:
$\tilde{V}(p\boldsymbol{a},px,p^{s},q,z)\equiv\tilde{V}(\boldsymbol{a},x,p^{s-1},q^{p},z^{p})\pmod{p^{s}}.$
Proof.
The normalized K-theoretic vertex function is
$\tilde{V}(p\boldsymbol{a},px,p^{s},q,z)=z^{px}\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{i=1}^{n}\dfrac{[q^{px-pa_{i}+d},q]_{p^{s}}}{[1,q]_{p^{s}}}\right)z^{d}$
Lemma 4.1 gives
$z^{px}\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{i=1}^{n}\dfrac{[q^{px-pa_{i}+d},q]_{p^{s}}}{[1,q]_{p^{s}}}\right)z^{d}\equiv z^{px}\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{i=1}^{n}\dfrac{[q^{px-pa_{i}+pd},q^{p}]_{p^{s-1}}}{[1,q^{p}]_{p^{s-1}}}\right)z^{pd}\pmod{p^{s}}$
which is exactly the statement of the theorem. ∎
Since the vertex describes all solutions of the $q$-holonomic system we also obtain:
@@
The normalized fundamental solution matrices satisfy the following congruences:
$\widetilde{\Psi}(p\boldsymbol{a},p^{s},q,z)\pmod{p^{s}}=\widetilde{\Psi}(\boldsymbol{a},p^{s-1},q^{p},z^{p}).$
4.3.
To prove Theorem 3.2 it is enough to show that for any $s\in\mathbb{N}$,
$\mathsf{U}(z)=\widetilde{\Psi}(p\boldsymbol{a},ph,q,z)\,\widetilde{\Psi}(\boldsymbol{a},h,q^{p},z^{p})^{-1}\pmod{p^{s}},$
is a rational function in $z$. Let us write $h=h_{0}+p^{s-1}h_{1}$ for some $h_{0}\in\mathbb{N}$ and $h_{i}\in\mathbb{Z}_{p}$. By Theorem 2.1 we have:
$\mathsf{U}(z)=\widetilde{\Psi}(p\boldsymbol{a},ph,q,z)=\widetilde{\Psi}(p\boldsymbol{a},ph_{0}+h_{1}p^{s},q,z)=\left(\prod\limits_{i=h_{1}p^{s}}^{ph_{0}+h_{1}p^{s}-1}S^{h}(p\boldsymbol{a},i,q,z)\right)\widetilde{\Psi}(p\boldsymbol{a},h_{1}p^{s},q,z),$
where $S^{h}(p\boldsymbol{a},i,q,z)$ is a certain invertible matrix whose coefficients are rational functions of $z$. By Corollary 4.3 we obtain:
$\widetilde{\Psi}(p\boldsymbol{a},ph,q,z)\equiv\left(\prod\limits_{i=0}^{ph_{0}-1}S^{h}(p\boldsymbol{a},i,q,z)\right)\widetilde{\Psi}(\boldsymbol{a},h_{1}p^{s-1},q^{p},z^{p})\pmod{p^{s}}.$
Similarly, we obtain
$\widetilde{\Psi}(\boldsymbol{a},h,q^{p},z^{p})\equiv\left(\prod\limits_{i=0}^{h_{0}-1}S^{h}(\boldsymbol{a},i,q^{p},z^{p})\right)\widetilde{\Psi}(\boldsymbol{a},h_{1}p^{s-1},q^{p},z^{p})$
and thus
(4.2) $\widetilde{\Psi}(p\boldsymbol{a},ph,q,z)\,\widetilde{\Psi}(\boldsymbol{a},h,q^{p},z^{p})^{-1}\equiv\left(\prod\limits_{i=0}^{ph_{0}-1}S^{h}(p\boldsymbol{a},i,q,z)\right)\left(\prod\limits_{i=0}^{h_{0}-1}S^{h}(\boldsymbol{a},i,q^{p},z^{p})\right)^{-1}\pmod{p^{s}}$
which is a product of finitely many matrices whose entries are rational functions in $z$.
4.4.
Finally, from the normalization (2.2), (2.3) we find:
$\mathsf{U}(0)=\dfrac{\Phi(p\boldsymbol{a},ph,px,q,z)}{\Phi(\boldsymbol{a},h,x,q^{p},z^{p})}$
From definition (2.2) we see that $\mathsf{U}(0)$ is exactly the product of factors as in Lemma 4.4 below, and therefore
$\mathsf{U}(0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p,q}(px+ph-pa_{j})}{\Gamma_{p,q}(px-pa_{j})\Gamma_{p,q}(ph)}$
which finishes the proof of the theorem.
4.5.
If $n$ and $h$ are two natural numbers then
$\dfrac{(q^{ph},q)_{\infty}}{(q,q)_{\infty}}\dfrac{(q^{pn+1},q)_{\infty}}{(q^{pn+ph},q)_{\infty}}=\binom{pn+ph-1}{pn}_{\!\!q}.$
is a $q$-binomial coefficient (3.3). Similarly
For the ratio of these factors we find:
where the last equality is by definition of $(p,q)$-Gamma function. By Koblitz’s Theorem 3.1 we obtain
Lemma 4.4.
The function of $n,h\in\mathbb{N}$ given by
$\dfrac{\dfrac{(q^{ph},q)_{\infty}}{(q,q)_{\infty}}\dfrac{(q^{pn+1},q)_{\infty}}{(q^{pn+ph},q)_{\infty}}}{\dfrac{(q^{ph},q^{p})_{\infty}}{(q^{p},q^{p})_{\infty}}\dfrac{(q^{pn+p},q^{p})_{\infty}}{(q^{pn+ph},q^{p})_{\infty}}}=\dfrac{\Gamma_{p,q}(pn+ph)}{\Gamma_{p,q}(pn)\,\Gamma_{p,q}(ph)}$
extends to $\mathbb{Z}_{p}$ values of arguments $n$ and $h$.
5. Numerical examples of Theorem 3.2
5.1.
Let ${\Psi}(\boldsymbol{a},\hbar,q,z)$ be the fundamental solution matrix (1.12) for some fixed parameters. As in (3.2) we look for the Frobenius intertwiner matrix in the form
(5.1) $\mathsf{U}(z)={\Psi}(p\boldsymbol{a},ph,q,z)\Lambda{\Psi}(\boldsymbol{a},h,q^{p},z^{p})^{-1}.$
where $\Lambda$ is a constant matrix. In order to be a Frobenius intertwiner, we need to find such $\Lambda\in\mathsf{GL}_{n}(\mathbb{K})$ for which (5.1) is an element of $\mathsf{GL}_{n}(\widehat{\mathsf{R}}_{p})$. Explicitly, $\mathsf{U}(z)$ must be a power series
$\mathsf{U}(z)=\mathsf{U}_{0}+\mathsf{U}_{1}z+\dots\in\mathsf{Mat}_{n}(\mathbb{Z})[[z]]$
such that for any natural number $s$, there exists a natural number $m_{s}$ for which
In words: after multiplication by sufficiently large power of $(1-z)$ the matrix elements of $\mathsf{U}(z)$ are polynomials in $z$ modulo $p^{s}$. This condition is very strong and can be used to effectively compute the Frobenius intertwiners as we about to show.
5.2.
Note that if the parameters $\boldsymbol{a}=(a_{1},\dots,a_{n})$ are pairwise distinct the matrix $\Lambda$ must be diagonal. Indeed, the fundamental solution matrix $\tilde{\Psi}(\boldsymbol{a},\hbar,q,z)$ contains logarithmic terms given by matrix $\mathsf{E}(z)$ in (1.13). The intertwiner (5.1) is a power series in $z$ only if these log terms are cancelled. It is elementary to check that for the pairwise distinct $a_{i}$ this happens only if $\Lambda$ is diagonal. Since $\Lambda$ is defined up to a scalar factor, we may assume that $\Lambda_{n,n}=1$ so
for some unknown $c_{1},\dots,c_{n-1}\in\mathbb{Z}_{p}$.
5.3.
Let us assume $n=2$. In this case
(5.11) $\Lambda=\left(\begin{array}[]{cc}c&0\\ 0&1\end{array}\right).$
Let us pick a prime $p=3$. To meet the condition (3.1) let us also set
$q=1+p=4,$
and pick some random values of $a_{1},a_{2},h\in\mathbb{Z}_{p}$, for instance
(5.12) $h=\dfrac{1}{2},\ \ a_{1}=\dfrac{1}{5},\ \ a_{2}=0.$
$\bar{h}=h\pmod{p^{7}}=1094,\ \ \bar{a_{1}}=a_{1}\pmod{p^{7}}=875,\ \ \ \bar{a_{2}}=0.$
$c=c^{(0)}+c^{(1)}3+c^{(2)}3^{2}+\dots$
where $c^{(i)}\in\{0,1,2\}$.
Step 1: determination of $c^{(0)}$ and $c^{(1)}.$
Using a computer we approximate the power series ${\Psi}(\boldsymbol{a},\hbar,q,z)$ by expansion up to $O(z^{20})$. Substituting this approximation to (5.1) and trying the integers $0,1,2$ for the values of $c^{(0)}$ and $c^{(1)}$ we find that
$\mathsf{U}(z)\pmod{p}$
exists (i.e. no powers of $p$ in the denominators) only if $c^{(0)}=1$ and $c^{(1)}=0$. In which case we find that
Step 2: determination of $c^{(2)}.$
If $c^{(2)}=0$, using a computer we find
If $c^{(2)}=1$ we find
If $c^{(2)}=2$ we find
Step 3: determination of $c^{(3)}.$
If $c^{(3)}=0$ we find:
If $c^{(3)}=1$ we find:
If $c^{(3)}=2$ we find:
Continuing this process we will find that at $m$-th step there is always unique value of $c^{m}$ for which rationality condition (5.2) is met. For instance, up to the order $9$ these method gives the following value
[…]
5.4.
Theorem 3.2 and Corollary 3.3 gives the following formula for the constant $c$:
at (5.12) we have
For $p=3$ and $q=1+p=4$ we compute:
[…]
The author have performed the same numerical experiments for numerous other rational values of parameters $a_{1},a_{2}$ and $h$ in $\mathbb{Z}_{p}$. In all these cases the results are in agreement with Theorem 3.2. Same method can be used for ranks $n>2$, in which case we have to determine the numerical values of $n-1$ constants in (5.8). Again, we find full agreement with Theorem 3.2.
6. Limit to cohomology
6.1.
As $q\to 1$, the quantum difference equation of $X$ degenerates to the quantum differential equation, which governs quantum cohomology of $X$. In this limit, the equivairnat K-theory ring (2.1) is replaced by the equivairnat cohomology
In this limit the gamma class (2.2) takes the form:
(6.2) $\Gamma(\boldsymbol{a},h,x,z)=z^{x}\,\prod\limits_{i=1}^{n}\dfrac{\Gamma(x+h-a_{i})}{\Gamma(x+1-a_{i})\Gamma(h)}\in\hat{H}^{\bullet}_{T}(X).$
where $\Gamma$ denotes the classical Gamma function, and $\hat{H}^{\bullet}_{T}(X)$ is the appropriate completion of the cohomology ring containing the gamma class. The normalized cohomological vertex function (also known as $J$-function in cohomology) is:
$\widetilde{V}(\boldsymbol{a},h,x,z)=\sum\limits_{d=0}^{\infty}\,\Gamma(\boldsymbol{a},h,x+d,z)\,\in\,\hat{H}^{\bullet}_{T}(X)[[z]].$
The vertex function satisfy the differential equation
(6.3) $P(\boldsymbol{a},h,z)\widetilde{V}(\boldsymbol{a},h,x,z)=(x-a_{1})\dots(x-a_{n})\Gamma(\boldsymbol{a},h,x,z)=0$
(note that the right side is zero in (6.1)), where
(6.4) $P(\boldsymbol{a},h,z)=z\prod\limits_{i=1}^{n}\,(D+h-a_{i})-\prod\limits_{i=1}^{n}(D-a_{i})$
and $D=z\frac{d}{dz}$. The components of the vertex in some basis of $H^{\bullet}_{T}(X)$ provide a basis of solutions for the hypergeometric differential equation (6.3). In particular, in the basis given by the classes of the $T$-fixed points $X^{T}$, we obtain the standard generazlized hypergeometric series:
$\widetilde{V}(\boldsymbol{a},h,a_{i},z)=\Gamma(\boldsymbol{a},h,a_{i},z)\,\sum\limits_{d=0}^{\infty}\,\left(\prod\limits_{j=1}^{n}\dfrac{(a_{i}+h-a_{j})_{d}}{(a_{i}+1-a_{j})_{d}}\right)z^{d}$
where $(a)_{d}=a(a+1)\dots(a+d-1)$.
By Theorem 3.1 $\Gamma_{p,q}(n)\stackrel{{\scriptstyle q\to 1}}{{\longrightarrow}}\Gamma_{p}(n)$ and from Theorem 3.2 we deduce:
Theorem 6.1.
For any $\boldsymbol{a}=(a_{1},\dots,a_{n})\in\mathbb{Z}^{n}_{p},\ \ h\in\mathbb{Z}_{p}$ there exists unique Frobenius intertwiner $\mathsf{U}(z)$ from the hypergeometric system (6.4) with parameters $(a_{1},\dots,a_{n},h,z^{p})$ to the system with parameters $(pa_{1},\dots,pa_{n},ph,z)$. The constant terms of the intertwiner is represented by the operator acting in $\hat{H}^{\bullet}_{T}(X)$ as an operator of multiplication by the cohomology class class
(6.5) $\mathsf{U}(0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p}(px+ph-pa_{j})}{\Gamma_{p}(px-pa_{j})\Gamma_{p}(ph)},$
where $\Gamma_{p}$ is the Morita gamma function.
The analog of Corollary 3.3 also gives:
Corollary 6.2.
If $a_{1},\dots,a_{n}$ are pairwise distinct, the constant term of the Frobenius intertwiner in the basis of torus fixed points is given by diagonal matrix with the following eigenvalues
(6.6) $\mathsf{U}(0)_{i,i}=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p}(p(a_{i}-a_{j}+h))}{\Gamma_{p}(p(a_{i}-a_{j}))\Gamma_{p}(ph)}.$
In this way we arrive at the description of the Frobenius intertwines for the $p$-adic hypergeometric equations discovered by Dwork [Dw89] see also [Ked21] where this result was recently revisited by other tools.
Remark.
Recall that for $n\in p\mathbb{Z}_{p}$ the Morita Gamma function satisfies the identities $\Gamma_{p}(n)=-\Gamma_{p}(1+n)=(-1)^{n_{0}}/\Gamma_{p}(-n)$. Using this, up to a scalar multiple, the constant term of Frobenius (6.5) takes the form
$\mathsf{U}(0)=\prod\limits_{i=1}^{n}\,\Gamma_{p}(px+ph-pa_{j})\Gamma_{p}(pa_{j}-px)$
The arguments of gamma functions in this product are exactly the $p$-multiples of the Chern roots of the tangent bundle $TX$. We thus can write this formally as
(6.7) $\mathsf{U}(0)=\Gamma_{p}(TX^{(p)}),$
where $(p)$ indicates the Frobenius twist multiplying the Chern roots by $p$. We expect that (6.7) holds for the constant term of the Frobenius intertwiner of the quantum differential equation for an arbitrary $X$.
@@ 7. Limit to $X=\mathbb{P}^{n-1}$: Frobenius structure for generalized Bessel equations
The quantum differential equation and its solutions for $X=\mathbb{P}^{n-1}$ can be obtained as the scaling limit $\hbar\to\infty$ of the corresponding equations and solutions for $T^{*}\mathbb{P}^{n-1}$ (we recall that the equivariant parameter $\hbar$ is the character corresponding to the dilation of the cotangent directions of $T^{*}\mathbb{P}^{n-1}$ under the torus action) [SV24]. In this section we consider the Frobenius intertwiners for the quantum differential equation of $\mathbb{P}^{n-1}$. We also describe the Frobenius automorphism for the non-equivariant limit of this equations which correspond to the generalized Bessel equations.
@@
$P(\boldsymbol{a},z)\widetilde{V}(x,\boldsymbol{a},z)=0.$
Formal limit $h\to\infty$ of Theorem 6.1 then gives:
@@
(7.3) $\mathsf{U}(0)=\prod\limits_{j=1}^{n}{\Gamma_{p}}(px-pa_{j}).$
We note that the arguments of the gamma functions in the product (7.3) are the $p$ multiples of the Chern roots of the tangent bundle $TX$. Thus, the universal formula (6.7) holds, in agreement with the Remark of the previous Section.
@@
The functions $\widetilde{V}_{0}(z),\dots,\widetilde{V}_{n}(z)$ provide a basis of solutions to the generalized Bessel equation. Note that (7.5) has a multiple $z^{x}$ which has the following expansion:
@@
(7.13) $(\widetilde{V}_{0}(z),\dots,\widetilde{V}_{n-1}(z))=({V}_{0}(z),\dots,{V}_{n-1}(z))\cdot\left(\begin{array}[]{ccccc}1&\ln(z)&\ln(z)^{2}/2&..&\ln(z)^{n-1}/(n-1)!\\ 0&1&\ln(z)&..&\ln(z)^{n-2}/(n-2)!\\ 0&0&1&..&\dots\\ ..&..&..&..&\dots\\ 0&0&0&..&1\end{array}\right)$
where $V_{i}(z)$ are some analytic (power series in $z$) functions. In particular, $\tilde{V}_{0}(z)$ does not contain logarithmic terms. For $n=2$, $\tilde{V}_{0}(z)$ is the power series representing the classical Bessel function of the first kind.
@@
and the exponent of a nilpotent operator is, of course, elementary to compute.
@@
The existence of the Frobenius automorphism for the Bessel equation was discovered by B. Dwork [Dwo74]. In particular, in Lemma 3.2 of [Dwo74] he gave an explicit formula for the matrix $\mathsf{U}(0)$ in terms of certain converging $p$-adic series. One checks that Dwork’s matrix $\mathfrak{U}(0)$ coincides with (7.18) after transposition. In particular, the series for the coefficients $\mathfrak{U}_{2,1}(0)=\gamma$ converges to $p\Gamma^{\prime}_{p}(0)$.
The action of the Frobenius automorphism for the generalized Bessel equations ($n>2$) was further studied by S. Sperber in [Sp80]. The results of this Section are in full agreement with these calculations too.
8. Frobenius intertwiners at $p^{s}$ roots of unity
@@
Let $\mathsf{U}(z)$ be the Frobenius intertwiner for the $q$-difference equation as in Theorem 3.2. As we expain in Section 3.2, the defining property of the Frobenius intertwiner $\mathsf{U}(z)$ discussed in the previous Sections is the condition
$\mathsf{U}(z)\pmod{p^{s}}={\mathrm{rational\ function\ of}}\ \ z$
In fact, a stronger property holds
$\mathsf{U}(z)\pmod{[p^{s}]_{q}}={\mathrm{rational\ function\ of}}\ \ z,$
where
$[p^{s}]_{q}=\dfrac{1-q^{p^{s}}}{1-q}=\varphi_{p}(q)\varphi_{p}(q^{p})\dots\varphi_{p}(q^{p^{s-1}})$
and $\varphi_{p}(q)=1+q+\dots+q^{p-1}$ is the $p$ th cyclotomic polynomial. This result follows from:
Theorem 8.1.
For $s>0$ let $\zeta\in\mathbb{K}$ denotes a primitive $p^{s}$ th roots of unity, then
$\left.\mathsf{U}(z)\right|_{q=\zeta}$
is well defined and is a rational function of $z$.
Proof.
At $q=\zeta$ we have the following (obvious) identity
$(z,q)_{p^{s}}=(z^{p},q^{p})_{p^{s-1}}=(1-z^{p^{s}})$
Arguing as in the proof of Lemma 4.1 we would find that at $q=\zeta$:
$\dfrac{[q^{i},q]_{p^{s}}}{[1,q]_{p^{s}}}=\left\{\begin{array}[]{ll}\dfrac{[q^{i},q^{p}]_{p^{s-1}}}{[1,q^{p}]_{p^{s-1}}}=1,&\mathrm{if}\ \ p\mid i,\\ &\\ 0,&\mathrm{if}\ \ p\nmid i.\end{array}\right.$
We then use this result as the analog of Lemma 4.1 to prove that the intertwiner (4.2) is rational, following the same steps as in the proof of Theorem 3.2. ∎
Remark.
For the case $n=1$ this result was recently obtained by D. Vargas-Montoya in [VM22]. In the terminology of [VM22] Theorem 8.1 can be formulated as existence of the strong Frobenius structure for the $q$-hypergeometric equation (1.2).

### v2

@@ 6. Limit to cohomology
For any $\boldsymbol{a}=(a_{1},\dots,a_{n})\in\mathbb{Z}^{n}_{p},\ \ h\in\mathbb{Z}_{p}$ the power series $\mathsf{U}(h,\boldsymbol{a},z)$ takes values in $\mathsf{GL}_{n}(E_{p})$ and gives a Frobenius intertwiner from the hypergeometric system (6.5) with parameters $(a_{1},\dots,a_{n},h,z^{p})$ to the hypergeometric system with parameters $(pa_{1},\dots,pa_{n},ph,z)$. The constant term of this intertwiner is the operator acting in $\hat{H}^{\bullet}_{T}(X)$ as an operator of multiplication by the cohomology class class
(6.6) $\mathsf{U}(h,\boldsymbol{a},0)=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p}(px+ph-pa_{j})}{\Gamma_{p}(px-pa_{j})\Gamma_{p}(ph)},$
where $\Gamma_{p}$ is the Morita gamma function.
The $q\to 1$ limit of Corollary 5.4 also gives:
@@
If $a_{1},\dots,a_{n}$ are pairwise distinct, then the constant term of the Frobenius intertwiner $\mathsf{U}(h,\boldsymbol{a},z)$ is diagonal in the basis of $H^{*}_{T}(X)$ given by the classes of $T$-fixed points with the eigenvalues
(6.7) $\mathsf{U}(h,\boldsymbol{a},0)_{i,i}=\prod\limits_{j=1}^{n}\,\dfrac{\Gamma_{p}(p(a_{i}-a_{j}+h))}{\Gamma_{p}(p(a_{i}-a_{j}))\Gamma_{p}(ph)},\ \ \ i=1,\dots,n.$
In this way we arrive at the description of the Frobenius intertwines for the $p$-adic hypergeometric equations discovered by Dwork [Dw69, Dw89] see also [Ked21] where this result was recently revisited by other tools.
Recall that for $n\in p\mathbb{Z}_{p}$ the Morita gamma function satisfies the identities $\Gamma_{p}(n)=-\Gamma_{p}(1+n)=1/\Gamma_{p}(-n)$. Using this, up to a constant scalar multiple independent of $x$, (6.6) can we written in the form
$\mathsf{U}(h,\boldsymbol{a},0)=\prod\limits_{i=1}^{n}\,\Gamma_{p}(px+ph-pa_{j})\Gamma_{p}(pa_{j}-px)$
The arguments of the gamma functions in this product are exactly the $p$-multiples of the Chern roots of the tangent bundle $TX$, i.e., they are the Chern roots of $TX^{(p)}:=\psi^{p}(TX)$ where $\psi^{p}$ is the $p$-th Adams operation. We thus can write this formally as
Proposition 6.3.
The constant term of the intertwiner $\mathsf{U}(h,\boldsymbol{a},{\boldsymbol{z}})$ acts in the equivariant cohomology $H^{*}_{T}(X)$ as multiplication by the cohomology class
(6.8) $\mathsf{U}(h,\boldsymbol{a},0)=\Gamma_{p}(TX^{(p)})$
Here, the cohomology class $\Gamma_{p}({\mathcal{V}})$ of a vector bundle $\mathcal{V}$ is defined as in (5.7) with $\Gamma_{p,q}$ replaced by the Morita gamma $\Gamma_{p}$. We expect that (6.8) holds for the constant term of the Frobenius intertwiner of the quantum differential equations associated with more general varieties than $X$. The last Proposition is the cohomological limit of Proposition 5.5.
@@ 7. Limit to $X=\mathbb{P}^{n-1}$: Frobenius structure for generalized Bessel equations
The equivariant quantum differential equation and its solutions for $X=\mathbb{P}^{n-1}$ arise as the limit $\hbar\to\infty$ of the corresponding equations and solutions for $T^{*}\mathbb{P}^{n-1}$ (we recall that the equivariant parameter $\hbar$ is the character of $T$ corresponding to the dilation of the cotangent directions of $T^{*}\mathbb{P}^{n-1}$ under the $T$-action) [SV24]. Thus, the formal limit $\hbar\to\infty$ of our results in the previous section gives the Frobenius intertwiners for the $T$-equivariant quantum differential equations of projective spaces.
Further, in the non-equivariant limit, when we ignore the $T$-action on $X=\mathbb{P}^{n-1}$, the quantum differential equation is nothing but the generalized Bessel equation (and the classical Bessel differential equation for $n=2$). The non-equivariant limit corresponds to the case $\boldsymbol{a}=(0,\dots,0)$. Specializing at this point, we arrive at the Frobenius structure for Bessel equations studied in [Dw74, Sp80]. This section is a summary of these limits.
@@
(7.3) $P(\boldsymbol{a},z)\widetilde{V}(x,\boldsymbol{a},z)=0.$
Let $\mathsf{U}(h,\boldsymbol{a},z)$ be as in Theorem 6.1 and let $\mathsf{U}(\boldsymbol{a},z)$ denotes its limit as $\hbar\to\infty$. Then we obtain
@@
(7.4) $\mathsf{U}(\boldsymbol{a},0)=\prod\limits_{j=1}^{n}{\Gamma_{p}}(px-pa_{j}).$
We note that the arguments of the gamma functions in the product (7.4) are the $p$ multiples of the Chern roots of the tangent bundle $TX$. Thus, the universal formula (6.8) still holds.
@@
The functions $\widetilde{V}_{0}(z),\dots,\widetilde{V}_{n}(z)$ provide a basis of solutions to the generalized Bessel equation. Note that (7.6) includes a multiple $z^{x}$ which has the following expansion:
@@
(7.14) $\ \ \ \ \ \ \ \ \tiny(\widetilde{V}_{0}(z),\dots,\widetilde{V}_{n-1}(z))=({V}_{0}(z),\dots,{V}_{n-1}(z))\cdot\left(\begin{array}[]{ccccc}1&\ln(z)&\ln(z)^{2}/2&..&\ln(z)^{n-1}/(n-1)!\\ 0&1&\ln(z)&..&\ln(z)^{n-2}/(n-2)!\\ 0&0&1&..&\dots\\ ..&..&..&..&\dots\\ 0&0&0&..&1\end{array}\right)$
where $V_{i}(z)$ are some power series in $z$. In particular, $\tilde{V}_{0}(z)$ does not contain logarithmic terms. For $n=2$, $\tilde{V}_{0}(z)$ is the power series representing the classical Bessel function of the first kind.
@@
and the exponent of a nilpotent operator is, of course, well defined.
@@
The existence of the Frobenius automorphism for the Bessel equation was discovered by Dwork [Dw74]. In particular, in Lemma 3.2 of [Dw74] he gave an explicit formula for the matrix $\mathsf{U}(0)$ in terms of certain converging $p$-adic series. One checks that Dwork’s matrix $\mathfrak{U}(0)$ coincides with (7.19) after transposition. In particular, his series for the coefficient $\mathfrak{U}_{2,1}(0)=\gamma$ converges to $p\Gamma^{\prime}_{p}(0)$.
The action of the Frobenius automorphism for the generalized Bessel equations ($n>2$) was further studied by Sperber in [Sp80]. The results of this Section are also in full agreement with these results.
We note that the values of $p$-adic zeta function appear naturally in our analysis of hypergeometric equations when we pass from the equivariant cohomology to the non-equivariant cohomology. Namely, the values of $p$-adic zeta functions appear as the coefficients of the expansion of the cohomology class $\Gamma_{p}(TX^{(p)})$ is some basis of $H^{*}(X)$. We expect the values of $p$-adic zeta appearing in the study of the mirror symmetry for more general examples of $X$, e.g., [BV23], should be treated in the same way.
8. Numerical examples of Theorem 5.2
@@
Let ${\Psi}(h,\boldsymbol{a},q,z)$ be the fundamental solution matrix (2.11). For simplicity we assume that all parameters are rational. As in (5.1) we look for the Frobenius intertwiner matrix in the form
(8.1) $\mathsf{U}(z)={\Psi}(ph,p\boldsymbol{a},q,z)\Lambda{\Psi}(h,\boldsymbol{a},q^{p},z^{p})^{-1}.$
where $\Lambda$ is a constant matrix. In order to be a Frobenius intertwiner, we need to find $\Lambda$ for which (8.1) is an element of $\mathsf{GL}_{n}(E_{p})$. Lemma 4.1 says that $\mathsf{U}(z)$ must be a power series
$\mathsf{U}(z)=\mathsf{U}_{0}+\mathsf{U}_{1}z+\dots\in\mathsf{Mat}_{n}(\mathbb{Q})[[z]]$
such that for any natural number $s$, there exists a natural number $m_{s}$ for whichLemma 4.1 says that $\mathsf{U}(z)\pmod{p^{s}}$ must be a rational function. Denominators of these rational functions are controlled by the singularities of the corresponding differential equations. In our case, the qde has a singularity at $z=1$ and therefore the denominators must be powers of $(1-z)$.
In words: after multiplication by sufficiently large power of $(1-z)$ the matrix elements of $\mathsf{U}(z)$ reduced $\pmod{p^{s}}$ truncate to polynomials in $z$. This condition is very strong and can be used to effectively discover non-trivial Frobenius intertwiners as we are about to show.
8.2.
Note that if the parameters $\boldsymbol{a}=(a_{1},\dots,a_{n})$ are pairwise distinct the matrix $\Lambda$ must be diagonal. Indeed, the fundamental solution matrix ${\Psi}(\boldsymbol{a},\hbar,q,z)$ contains logarithmic terms given by matrix $\mathsf{E}(z)$ in (2.12). The intertwiner (8.1) is a power series in $z$ only if these log terms are canceled. It is elementary to check that for the pairwise distinct $a_{i}$ this happens only if $\Lambda$ is diagonal. Since $\Lambda$ is defined up to a scalar factor, we may assume further that $\Lambda_{n,n}=1$ so
for some unknown $c_{1},\dots,c_{n-1}\in\mathbb{Z}_{p}$.
8.3.
Let us assume $n=2$. In this case
(8.11) $\Lambda=\left(\begin{array}[]{cc}c&0\\ 0&1\end{array}\right).$
Let us pick a prime $p=3$, set $q=1+p=4$, and pick some random values of $a_{1},a_{2},h\in\mathbb{Z}_{p}$, for instance
(8.12) $h=\dfrac{1}{2},\ \ a_{1}=\dfrac{1}{5},\ \ a_{2}=0.$
$\bar{h}=h\pmod{3^{7}}=1094,\ \ \bar{a_{1}}=a_{1}\pmod{3^{7}}=875,\ \ \ \bar{a_{2}}=0.$
$c=c^{(0)}+c^{(1)}3+c^{(2)}3^{2}+\dots$
where $c^{(i)}\in\{0,1,2\}$.
Step 1: determination of $c^{(0)}$ and $c^{(1)}.$
Using a computer we approximate the power series ${\Psi}(\boldsymbol{a},\hbar,q,z)$ by series expansion (2.11) up to $O(z^{20})$. Substituting this approximation to (8.1) and trying the integers $0,1,2$ for the values of $c^{(0)}$ and $c^{(1)}$ we find that $\mathsf{U}(z)\pmod{p}$ exists (i.e. no powers of $p$ in the denominators) only if $c^{(0)}=1$ and $c^{(1)}=0$, in which case we find that
Step 2: determination of $c^{(2)}.$
If $c^{(2)}=0$, using a computer we find
If $c^{(2)}=1$ we find
If $c^{(2)}=2$ we find
Step 3: determination of $c^{(3)}.$
If $c^{(3)}=0$ we find:
If $c^{(3)}=1$ we find:
If $c^{(3)}=2$ we find:
Continuing this process we will find that at $m$-th step there is always unique value of $c^{(m)}$ for which rationality condition (8.2) is met. Up to the order $9$ these method gives the following value
[…]
8.4.
Theorem 5.2 and Corollary 5.4 give the following formula for the constant $c$:
at (8.12) we have
For $p=3$ and $q=1+p=4$ we compute:
[…]
The author have performed the same numerical experiments for numerous other rational values of parameters $a_{1},a_{2}$ and $h$ in $\mathbb{Z}_{p}$. In all these cases the results are in full agreement with Theorem 5.2. Same method can be used for ranks $n>2$, in which case we have to determine the numerical values of $n-1$ constants in (8.8). Again, we find full agreement with Theorem 5.2.